OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1/4x + 2. y = 4x - 1 y = 4x - 3 y = 1/4x - 3 y = 1/4x - 1

OpenStudy (anonymous):

The slope will stay the same 1/4 y = 1/4x + b so find the value of b and put it into y = 1/4x + b

OpenStudy (anonymous):

-2 = 1/4(-4) + b solve for b

OpenStudy (anonymous):

Is it the last one?

OpenStudy (anonymous):

y+2=1/4(x+4) y=1/4x+1-2 y=1/4x -1

OpenStudy (anonymous):

I forgot about point-slope form...wow:/

OpenStudy (anonymous):

Nice job @myko :D